OpenStudy (anonymous):

Please assist me? I'll answer anything related to English, Psychology, or Philosophy in return? ;) Solve 2x + 5y = 10 for y. a.) y = 2x + 2 b.) y = -2x + 10 c.) y = x + 2 d.) y = x + 10

OpenStudy (anonymous):

none of the above

OpenStudy (anonymous):

there has to be one

OpenStudy (hba):

5y=10-2x ,y=10-2x/5

OpenStudy (anonymous):

y = (10-2x)/5

OpenStudy (hba):

Which means none of the above got it @InsanelyChaotic

OpenStudy (anonymous):

y = 2x + 2 y = -2x + 10 y= 2/5 x + 2 y= 2/5 x +10

OpenStudy (hba):

y=2-2/5 x

OpenStudy (hba):

Still none of the above

OpenStudy (anonymous):

recheck your question you are missing something

OpenStudy (anonymous):

Tell that to my professor, he'll say otherwise

OpenStudy (anonymous):

I will, one moment, my apologies

OpenStudy (hba):

YOU HAVE COPIED THE QUESTION WRONG

OpenStudy (anonymous):

y=2-2x/5

OpenStudy (anonymous):

Calm down, i'm revising

OpenStudy (hba):

oK

OpenStudy (anonymous):

a. )y = 2x + 2 b.) y = -2x + 10 c.) y= 2/5 x + 2 d.) y= 2/5 x +10

OpenStudy (hba):

What about the question ?

OpenStudy (anonymous):

Solve 2x + 5y = 10 for y.

OpenStudy (hba):

y=10-2/5x

OpenStudy (anonymous):

y = 2/5 x + 10 y = 2x + 2 y = - 2/5 x + 2 y = –2x + 10

OpenStudy (anonymous):

voila

OpenStudy (anonymous):

Which ones accurate.....?

OpenStudy (anonymous):

If I want to tell my professor off correctly, it needs to be with the right answer.

OpenStudy (anonymous):

y = -2/5 x + 2

OpenStudy (anonymous):

Otherwise, he'll go on and on about how poor at math I am

OpenStudy (anonymous):

Which is TRUE, but unnecessary

OpenStudy (anonymous):

So it's either B or C

OpenStudy (hba):

Well I am Sorry To Say But None Of The Options You Gave Match With The Answer :(

OpenStudy (anonymous):

On my paper I have C

OpenStudy (hba):

You need To concentrate in Class @InsanelyChaotic because i think you have problem copying the right question which results in getting the right answer :(

OpenStudy (hba):

Wrong answer *